# FC-Attack: Jailbreaking Large Vision-Language Models via Auto-Generated Flowcharts

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have become powerful and widely adopted in some practical applications. However, recent research has revealed their vulnerability to multimodal jailbreak attacks, whereby the model can be induced to generate harmful content, leading to safety risks. Although most LVLMs have undergone safety alignment, recent research shows that the visual modality is still vulnerable to jailbreak attacks.
In our work, we discover that by using flowcharts with partially harmful information, LVLMs can be induced to provide additional harmful details. Based on this, we propose a jailbreak attack method based on auto-generated flowcharts, $\mathsf{FC\text{-}Attack}$. Specifically, $\mathsf{FC\text{-}Attack}$ first fine-tunes a pre-trained LLM to create a step-description generator based on benign datasets. The generator is then used to produce step descriptions corresponding to a harmful query, which are transformed into flowcharts in $3$ different shapes (vertical, horizontal, and S-shaped) as visual prompts. These flowcharts are then combined with a benign textual prompt to execute a jailbreak attack on LVLMs. Our evaluations using the Advbench dataset show that $\mathsf{FC\text{-}Attack}$ achieves over $90\%$ attack success rates on Gemini-1.5, Llaval-Next, Qwen2-VL, and InternVL-2.5 models, outperforming existing LVLM jailbreak methods. Additionally, we investigate factors affecting the attack performance, including the number of steps and the font styles in the flowcharts. Our evaluation shows that $\mathsf{FC\text{-}Attack}$ can improve the jailbreak performance from $4\%$ to $28\%$ in Claude-3.5 by changing the font style.
To mitigate the attack, we explore several defenses and find that AdaShield can largely reduce the jailbreak performance but with the cost of utility drop.
@@
Figure 1: Comparison of jailbreak effectiveness in various LVLMs using three prompt types: (1) harmful text without images; (2) harmful step descriptions without images; and (3) harmful step descriptions in images.
@@ Introduction
With the advancement of Large Language Models (LLMs), Large Vision-Language Models (LVLMs) that integrate vision and text, such as GPT-4o [14], have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding [20, 16].
@@
Despite being powerful, recent studies [11, 27] have revealed that LVLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards in LVLMs and gain harmful knowledge. Such vulnerabilities pose remarkable safety risks to the Internet and the physical world. For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion. [^1] https://www.thetimes.com/world/us-world/article/vegas-cybertruck-bomber-used-chatgpt-to-plan-explosion-wk7h2jtkq. To better safeguard LVLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al. [42] providing a quantitative understanding regarding the adversarial vulnerability of LVLMs. Previous studies often create adversarial datasets tailored to specific models, which tend to perform poorly on other models.
Currently, jailbreak attacks against LVLMs can be broadly categorized into two main types: optimization-based attacks [5, 19] and prompt-based attacks [11, 33]. Optimization-based attacks involve identifying an image that semantically corresponds to a harmful textual prompt. Adversarial perturbations are then added to this image using gradient-based optimization techniques to achieve the jailbreak. These methods usually require white-box access and are relatively slow due to the optimization process. In contrast, prompt-based jailbreak methods require only black-box access, which can only observe the model’s outputs without direct access to its internal parameters. These attacks exploit the vulnerabilities in LVLMs’ alignment mechanisms, which focus mainly on text-based safety alignment. By injecting malicious information into the visual prompt, these methods guide LVLMs in generating harmful content. Generally, optimization-based methods are more effective in white-box scenarios but show limited transferability to black-box models. Prompt-based attacks are relatively straightforward. For example, FigStep merely embeds harmful queries into images without employing additional attack vectors, facilitating simpler defense against these attacks.
@@
To better improve the attack transferability and its effectiveness, we propose a novel prompt-based jailbreak attack, namely $\mathsf{FC\text{-}Attack}$. Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards. Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and Low-Rank Adaptation (LoRA) is applied to fine-tune a pre-trained LLM to obtain a step-description generator. (2) Jailbreak Deployment: This stage uses the generator mentioned above to produce execution steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts. Together with the benign textual prompt, the visual prompt will be fed into LVLM to achieve the jailbreak. Note that the harmful flowcharts are generated automatically without hand-crafted effort, which enhances their efficiency.
@@
Our evaluation on the Advbench dataset shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5. Although the ASR is lower on GPT-4o mini, GPT-4o, and Claude-3.5, we how later that it can be improved in certain ways. To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of LVLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure 1), numbers of descriptions, and font styles in flowcharts. These experiments show that LVLMs exhibit higher safety in the text modality but weaker in the visual modality. Moreover, we find that even flowcharts with a one-step harmful description can achieve high ASR, as evidenced by the Gemini-1.5 model, where the ASR reaches $86\%$. Furthermore, font styles in flowcharts also contribute to the ASR increase. For instance, when the font style is changed from “Times New Roman” to “Pacifico”, the ASR increases from $4\%$ to $28\%$ on the model with the lowest ASR (Claude-3.5) under the original style. To mitigate the attack, we consider several popular defense approaches, including Llama-Guard-3-11B-Vision [24], JailGuard [39], AdaShield-S [32], and AdaShield-A [32]. Among them, AdaShield-A demonstrates the best defense performance by reducing the average ASR from $58.6\%$ to $1.7\%$. However, it also reduces LVLM’s utility on benign datasets, which calls for more effective defenses.
Overall, our contributions are summarized as follows:
  • We propose a novel method, $\mathsf{FC\text{-}Attack}$ that jailbreaks LVLMs using auto-generated harmful flowcharts. Experiments conducted on Advbench show that $\mathsf{FC\text{-}Attack}$ achieves a higher ASR across multiple models compared to other LVLM jailbreak attacks.
@@
  • Our ablation study investigates the impact of different types of user queries, the number of steps, and the font style in flowcharts. We find that the font style could serve as a key factor to further improve the ASR, especially for safer LVLMs, revealing a novel attack channel in LVLMs.
@@
Large Vision-Language Model
@@
In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities [41, 7], which have driven the emergence of LVLMs [38] (also known as Multimodal Large Language Models, MLLMs). LVLMs combine visual understanding with language comprehension, showing promising capabilities in visual downstream tasks, including Visual Question Answering (VQA) [3, 17, 28], image captioning [13, 18], and visual commonsense reasoning [37, 30].
@@
In this paper, we consider both popular open-source and widely used production LVLMs. These LVLMs are the most widely used, and all of them have been aligned to ensure safety. Detailed information are introduced in Appendix A
@@
Figure 2: Overview of the $\mathsf{FC\text{-}Attack}$ framework, including two stages: Step-Description Generator Building and Jailbreak Deployment.
Jailbreak Attacks on LVLMs
Similar to LLMs, despite undergoing safety alignment, LVLMs remain susceptible to jailbreak attacks. Current attacks can be categorized into two types: optimization-based and prompt-based attacks (also known as non-optimization-based attacks).
@@
Most existing optimization-based attacks rely on backpropagating the gradient of the target to generate harmful outputs. These methods typically require white-box access to the model, where they obtain the output logits of LVLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts  [4, 29, 26] (e.g., the target can be “Sure! I’m ready to answer your question.”). Carlini et al. [6] are the first to propose optimizing input images by using fixed toxic outputs as targets, thereby forcing the model to produce harmful outputs. Building on this, Bailey et al. [5] introduce the Behaviour Matching Algorithm, which trains adversarial images to make LVLMs output behavior that matches a target in specific contextual inputs. This process requires the model’s output logits to align closely with those of the target behavior. Additionally, they propose Prompt Matching, where images are used to induce the model to respond to specific prompts. Li et al. [19] take this further by replacing harmful keywords in the original textual inputs with objects or actions in the image, allowing harmful information to be conveyed through images to achieve jailbreaking. Unlike previous work, these images are generated using diffusion models and are iteratively optimized with models like GPT-4. This approach enhances the harmfulness of the images, enabling more effective attacks.
@@
Unlike optimization-based attacks, prompt-based attacks only need black-box access to successfully attack the model without introducing adversarial perturbations into images. Gong et al. [11] discovers that introducing visual modules may cause the original security mechanisms of LLMs to fail in covering newly added visual content, resulting in potential security vulnerabilities. To address this, they propose the FigStep attack, which converts harmful textual instructions into text embedded in images and uses a neutral textual prompt to guide the model into generating harmful content. This method can effectively attack LVLMs without requiring any training. Wang et al. [33] identifies a phenomenon called Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs. Specifically, even if harmful instructions in text or images are rearranged, LVLMs can still correctly interpret their meaning. However, the safety mechanisms of LVLMs are often more easily bypassed by shuffled harmful inputs than by unshuffled ones, leading to dangerous outputs. Compared to optimization-based attacks, prompt-based attacks are simple to achieve high success rates against closed-source models. Our proposed $\mathsf{FC\text{-}Attack}$ also belongs to this category, requiring only black-box access to successfully jailbreak the model.
@@
Adversary’s Goal
The adversary’s goal is to exploit attack methods to bypass the protective mechanisms of LVLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy. [^2] https://openai.com/policies/usage-policies This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of powerful LVLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort. These objectives pose severe societal impacts and risks, turning model providers into indirect accomplices, and thereby damaging their reputations.
Adversary’s Capabilities
In this paper, we consider a black-box scenario where the adversary cannot directly access the model’s structure, parameters, or output logits, but can only obtain the model’s final output (texts). In this scenario, adversaries interact with the model through an API provided by the model owner. Moreover, the interaction is limited to a single-turn conversation, with no history stored beyond the predefined system prompt. This scenario is common in real-world applications, as many powerful models are closed-source, like GPT-4o, or adversaries lack the resources to deploy open-source models. Consequently, they can only access static remote instances via APIs while maintaining some flexibility in selecting different generation parameters such as temperature.
@@
Dataset Construction. To construct the Step-Description Dataset, we randomly select a topic $t\in\mathcal{T}$ from a collection of ordinary daily topics $\mathcal{T}$. Based on the selected topic $t$, we design a set of few-shot examples $\mathcal{S}$ and combine them into a complete prompt $P=\text{Compose}(t,\mathcal{S})$. This prompt is then fed into an LLM (gpt-4o-2024-08-06 in our evaluation) to generate action statements and step-by-step descriptions related to topic $t$, as shown below:
@@
The Generator $\mathcal{G}$ is capable of breaking down a task (query) into a series of detailed step descriptions based on the query. Given a query $q$ about the steps, $\mathcal{G}(q)$ represents the step-by-step solution given by the generator, where we find that it can also generate step descriptions for harmful queries after fine-tuning.
@@
After obtaining the Step-Description Generator $\mathcal{G}$, a harmful query $q_{h}$ is input to generate the corresponding step-by-step description. This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz [^3] https://graphviz.org.). Together with a benign textual prompt $p_{b}$ (more details are in Appendix B), the flowchart will be fed into the aligned LVLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
$o_{h}=\mathcal{A}(\mathcal{F}(\mathcal{G}(q_{h})),p_{b})\leftarrow\text{FC-attack}(q_{h}),$ (4)
@@
Target Model. We test $\mathsf{FC\text{-}Attack}$ on seven popular LVLMs, including the open-source models Llava-Next ( llama3-llava-next-8b) [22], Qwen2-VL (Qwen2-VL-7B-Instruct) [31], and InternVL-2.5 (InternVL-2.5-8B) [9] as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) [25], GPT-4o (gpt-4o-2024-08-06) [14], Claude (claude-3-5-sonnet-20240620) [2], and Gemini (gemini-1.5-flash) [12].
@@
Table 1: Comparison of ASR performance across different methods and LVLMs. (“Ensemble” in this paper is defined as a no-attack harmful query being considered successfully jailbroken if any of the three types of harmful flowcharts associated with it succeed in the jailbreak.)
@@
Evaluation Metric. In the experiments, we use the Attack Success Rate (ASR) to evaluate the performance of our attack, which can be defined as follows:
@@
Following the instructions of judge prompt [8, 34], we employ GPT-4o as the judge to evaluate whether the output is jailbroken or not.
FC-Attack Deployment. Referring to Section 4, $\mathsf{FC\text{-}Attack}$ consists of two stages. For the Step-Description Generator Building, we first use GPT-4o to randomly generate several daily topics and $3$ few-shot examples, which are then combined into a prompt and fed into GPT-4o to construct the dataset $\mathcal{D}_{t}$. In our experiments, the number of descriptions in the flowchart is limited to a maximum of $10$ steps, as too many descriptions can result in excessive length in one direction of the image. The dataset contains $5,000$ pairs of queries and step descriptions for daily activities, with the temperature set to 1 (more details are provided in Appendix C). We then select Mistral-7B-Instruct-v0.1 [15] as the pre-trained LLM and fine-tune it on $\mathcal{D}_{t}$ using LoRA. The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, $2$ epochs, a batch size of $8$, a learning rate of $51e-$ 5, and a weight decay of $51e-$ 5. For the jailbreak deployment stage, we set the temperature to $0.3$ for all LVLMs for a fair comparison.
Baselines. To better justify the effectiveness of $\mathsf{FC\text{-}Attack}$, we consider the following 5 jailbreak attacks against LVLMs as the baselines, which are categorized into black-box attacks (MM-SafetyBench [23], SI-Attack [40], and FigStep [11]) and white-box attacks (HADES [19], VA-Jailbreak [26]).
@@
Figure 3: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
@@
Figure 4: ASR under different prompts against LVLMs.
@@
In this section, we explore the performance of $\mathsf{FC\text{-}Attack}$ and conduct ablation study and defense research. We conduct jailbreak experiments on LVLMs for $\mathsf{FC\text{-}Attack}$. As shown in Figure 3 , it is a successful jailbreak case on Gemini-1.5.
Figure 5: Different styles of fonts in flowcharts (“1 step”).
Table 2: Comparison of ASR performance for different models at various description numbers. (Vertical: Top-to-bottom flow; Horizontal: Left-to-right flow; S-shaped: S-shaped path flow.)
@@
In Table 1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models. We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both open-source and production models compared to all baselines. For example, the ASRs are $94\%$, $92\%$, $90\%$, and $90\%$ on Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, respectively.
However, the ASR on some production models, such as Claude-3.5, GPT-4o, and GPT-4o mini, is relatively low, at $4\%$, $30\%$, and $10\%$, respectively. This might be because these production models have more advanced and updated visual safety alignment strategies.

### v2

# FC-Attack: Jailbreaking Multimodal Large Language Models via Auto-Generated Flowcharts

@@ Abstract
Multimodal Large Language Models (MLLMs) have become powerful and widely adopted in some practical applications. However, recent research has revealed their vulnerability to multimodal jailbreak attacks, whereby the model can be induced to generate harmful content, leading to safety risks. Although most MLLMs have undergone safety alignment, recent research shows that the visual modality is still vulnerable to jailbreak attacks.
In our work, we discover that by using flowcharts with partially harmful information, MLLMs can be induced to provide additional harmful details. Based on this, we propose a jailbreak attack method based on auto-generated flowcharts, $\mathsf{FC\text{-}Attack}$. Specifically, $\mathsf{FC\text{-}Attack}$ first fine-tunes a pre-trained LLM to create a step-description generator based on benign datasets. The generator is then used to produce step descriptions corresponding to a harmful query, which are transformed into flowcharts in $3$ different shapes (vertical, horizontal, and S-shaped) as visual prompts. These flowcharts are then combined with a benign textual prompt to execute the jailbreak attack on MLLMs. Our evaluations on Advbench show that $\mathsf{FC\text{-}Attack}$ attains an attack success rate of up to 96% via images and up to 78% via videos across multiple MLLMs. Additionally, we investigate factors affecting the attack performance, including the number of steps and the font styles in the flowcharts. We also find that $\mathsf{FC\text{-}Attack}$ can improve the jailbreak performance from $4\%$ to $28\%$ in Claude-3.5 by changing the font style. To mitigate the attack, we explore several defenses and find that AdaShield can largely reduce the jailbreak performance but with the cost of utility drop.
@@ Introduction
With the advancement of Large Language Models (LLMs), Multimodal Large Language Models (MLLMs) that integrate vision (images and videos) and text, such as GPT-4o [18] and Qwen2.5-VL [6], have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding [24, 20].
Figure 1: Comparison of jailbreak effectiveness in various MLLMs using three prompt types.
@@
Despite being powerful, recent studies [14, 32] have revealed that MLLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards and gain harmful knowledge. Such vulnerabilities pose remarkable safety risks to the Internet and the physical world. For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion [36]. To better safeguard MLLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al. [52] providing a quantitative understanding regarding the adversarial vulnerability of MLLMs. Previous studies often create adversarial datasets tailored to specific models, which tend to perform poorly on other models.
Currently, jailbreak attacks against MLLMs can be broadly categorized into two main types: optimization-based attacks [8, 23] and prompt-based attacks [14, 39]. Optimization-based attacks use white-box gradient methods to craft adversarial perturbations on visual prompt aligned with harmful text. They are effective but slow and have limited transferability in black-box scenarios. In contrast, prompt-based jailbreaks require only black-box access and work by injecting malicious visual cues into benign prompts to exploit MLLMs’ text-focused safety alignment.
@@
To better improve the attack transferability and its effectiveness, we propose a novel prompt-based jailbreak attack, namely $\mathsf{FC\text{-}Attack}$. Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts (images and videos) as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards. Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and fine-tune a pre-trained LLM to obtain a step-description generator. (2) Jailbreak Deployment: This stage uses the generator to produce steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts. Together with the benign textual prompt, the visual prompt is fed into MLLMs to achieve the jailbreak. Note that the harmful flowcharts are generated automatically without hand-crafted effort.
@@
Our evaluation on the Advbench dataset shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5. Although the ASR is lower on GPT-4o mini, GPT-4o, and Claude-3.5, we how later that it can be improved in certain ways. To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of MLLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure 1), numbers of descriptions, and font styles in flowcharts. These experiments show that MLLMs exhibit higher safety in the text modality but weaker in the visual modality. Moreover, we find that even flowcharts with a one-step harmful description can achieve high ASR, as evidenced by the Gemini-1.5 model, where the ASR reaches $86\%$. Furthermore, font styles in flowcharts also contribute to the ASR increase. For instance, when the font style is changed from “Times New Roman” to “Pacifico”, the ASR increases from $4\%$ to $28\%$ on the model with the lowest ASR (Claude-3.5) under the original style. To mitigate the attack, we consider several popular defense approaches, including Llama-Guard-3-11B-Vision [28], JailGuard [47], AdaShield-S [38], and AdaShield-A [38]. Among them, AdaShield-A demonstrates the best defense performance by reducing the average ASR from $58.6\%$ to $1.7\%$. However, it also reduces MLLM’s utility on benign datasets, which calls for more effective defenses.
Overall, our contributions are as follows:
  • In this work, we develop $\mathsf{FC\text{-}Attack}$, which leverages auto-generated harmful flowcharts to jailbreak MLLMs via both image and video modalities. To the best of our knowledge, this is the first approach to exploit the video modality for MLLM jailbreak.
@@
  • Experiments on Advbench demonstrate that $\mathsf{FC\text{-}Attack}$ consistently achieves better ASR across multiple models compared to existing MLLM jailbreak attacks. Our ablation study investigates the impact of different types of user queries, the number of steps, and the font style in flowcharts. We find that the font style could serve as a key factor to further improve the ASR, especially for safer MLLMs, revealing a novel attack channel in MLLMs.
@@
Multimodal Large Language Models
@@
In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities [51, 10], which have driven the emergence of MLLMs [46] (also known as Large Vision Language Models, LVLMs). MLLMs combine visual understanding with language comprehension, showing promising capabilities in visual downstream tasks, including Visual Question Answering (VQA) [3, 21, 33], image captioning [17, 22], and visual commonsense reasoning [45, 35]. Notably, some MLLMs are capable of processing both image and video inputs, enabling broader applications across multimodal scenarios.
@@
In this paper, we consider both popular open-source and widely used production MLLMs. These MLLMs are the most widely used, and all of them have been aligned to ensure safety. Detailed information are introduced in Appendix A.
@@
Figure 2: Overview of the $\mathsf{FC\text{-}Attack}$ framework with two stages.
Jailbreak Attacks on MLLMs
@@
Similar to LLMs, which have been shown to be vulnerable to jailbreak attacks [42], MLLMs also remain susceptible despite safety alignment. Current attacks can be categorized into two types: optimization-based and prompt-based attacks. Most existing optimization-based attacks rely on backpropagating the gradient of the target to generate harmful outputs. These methods typically require white-box access to the model, where they obtain the output logits of MLLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts [4, 34, 31] (e.g., the target can be “Sure! I’m ready to answer your question.”). Carlini et al. [9] are the first to propose optimizing input images by using fixed toxic outputs as targets, thereby forcing the model to produce harmful outputs. Building on this, Bailey et al. [8] introduce the Behaviour Matching Algorithm, which trains adversarial images to make MLLMs output behavior that matches a target in specific contextual inputs. This process requires the model’s output logits to align closely with those of the target behavior. Additionally, they propose Prompt Matching, where images are used to induce the model to respond to specific prompts. Li et al. [23] take this further by replacing harmful keywords in the original textual inputs with objects or actions in the image, allowing harmful information to be conveyed through images to achieve jailbreaking. Unlike previous work, these images are generated using diffusion models and are iteratively optimized with models like GPT-4. This approach enhances the harmfulness of the images, enabling more effective attacks.
@@
Unlike optimization-based attacks, prompt-based attacks only need black-box access to successfully attack the model without introducing adversarial perturbations into images. Gong et al. [14] discovers that introducing visual modules may cause the original security mechanisms of LLMs to fail in covering newly added visual content, resulting in potential security vulnerabilities. To address this, they propose the FigStep attack, which converts harmful textual instructions into text embedded in images and uses a neutral textual prompt to guide the model into generating harmful content. This method can effectively attack MLLMs without requiring any training. Wang et al. [39] identifies a phenomenon named Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs. Specifically, even if harmful instructions in text or images are rearranged, MLLMs can still correctly interpret their meaning. However, the safety mechanisms of MLLMs are often more easily bypassed by shuffled harmful inputs than by unshuffled ones, leading to dangerous outputs. Compared to optimization-based attacks, prompt-based attacks usually achieve higher success rates against closed-source models. Our proposed $\mathsf{FC\text{-}Attack}$ also belongs to this category, requiring only black-box access.
@@
Adversary’s Goal. The adversary’s goal is to exploit attacks to bypass the protective mechanisms of MLLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy [30]. This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of MLLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort. These objectives pose severe societal impacts and risks to the model providers.
Adversary’s Capabilities. In this paper, we consider a black-box scenario where the adversary cannot directly access the model’s structure, parameters, or output logits, but can only obtain the model’s final output (texts). In this scenario, adversaries interact with the model through an API provided by the model owner. Moreover, the interaction is limited to a single-turn conversation, with no history stored beyond the predefined system prompt. This scenario is common in real-world applications, as many powerful models are closed-source, like GPT-4o, or adversaries lack the resources to deploy open-source models. Consequently, they can only access static remote instances via APIs.
@@
Dataset Construction. To construct the Step-Description Dataset, we randomly select a topic $t\in\mathcal{T}$ from a collection of ordinary daily topics $\mathcal{T}$. Based on it, we design a set of few-shot examples $\mathcal{S}$ and combine them into a complete prompt $P=\text{Compose}(t,\mathcal{S})$. This prompt is then fed into an LLM (gpt-4o-2024-08-06 in our evaluation) to generate action statements and step-by-step descriptions related to topic $t$, as shown below:
@@
The Generator $\mathcal{G}$ is capable of breaking down a task (query) into a series of detailed step descriptions based on the query. Given a query $q$ about the steps, $\mathcal{G}(q)$ represents the step-by-step solution given by the generator, where we find that is can also generate step descriptions for harmful queries after fine-tuning.
@@
After obtaining the Step-Description Generator $\mathcal{G}$, a harmful query $q_{h}$ is input to generate the corresponding step-by-step description. This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz [16]). Together with a benign textual prompt $p_{b}$ (more details are in Appendix B), the flowchart will be fed into the aligned MLLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
$o_{h}=\mathcal{A}(\mathcal{F}(\mathcal{G}(q_{h})),p_{b})\leftarrow\text{FC-attack}(q_{h}).$ (4)
@@
Target Model. We test $\mathsf{FC\text{-}Attack}$ on seven popular MLLMs, including the open-source models Llava-Next ( llama3-llava-next-8b) [26], Qwen2-VL (Qwen2-VL-7B-Instruct) [37], and InternVL-2.5 (InternVL-2.5-8B) [12] as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) [29], GPT-4o (gpt-4o-2024-08-06) [18], Claude (claude-3-5-sonnet-20240620) [2], and Gemini (gemini-1.5-flash) [15]. Moreover, we also test $\mathsf{FC\text{-}Attack}$ on LLMs via video, including Qwen-vl-max (Qwen-vl-max-latest) [7],Qwen2.5-Omni [40] and LLaVA-Video-7B-Qwen2 [49].
@@
Table 1: Comparison of ASR performance across different methods and MLLMs. (“Ensemble” in this paper is defined as a no-attack harmful query being considered successfully jailbroken if any of the three types of harmful flowcharts associated with it succeed in the jailbreak.)
@@
Evaluation Metric. In the experiments, we use the ASR to evaluate the performance of our attack, which can be defined as follows:
@@
Following the judge prompt [11], we employ GPT-4o to serve as the evaluator.
FC-Attack Deployment. Referring to Section 4, $\mathsf{FC\text{-}Attack}$ consists of two stages. For the Step-Description Generator Building, we first use GPT-4o to randomly generate several daily topics and $3$ few-shot examples, which are then combined into a prompt and fed into GPT-4o to construct the dataset $\mathcal{D}_{t}$. In our experiments, the number of descriptions in the flowchart is limited to a maximum of $10$ steps, as too many descriptions can result in excessive length in one direction of the image. The dataset contains $5,000$ pairs of queries and step descriptions for daily activities, with the temperature set to 1 (more details are provided in Appendix C). We then select Mistral-7B-Instruct-v0.1 [19] as the pre-trained LLM and fine-tune it on $\mathcal{D}_{t}$ using LoRA. The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, $2$ epochs, a batch size of $8$, a learning rate of $51e-$ 5, and a weight decay of $51e-$ 5. For the jailbreak deployment stage, we set the temperature to $0.3$ for all MLLMs for a fair comparison.
Baselines. To validate the effectiveness of $\mathsf{FC\text{-}Attack}$, we adopt five jailbreak attacks as baselines, which are categorized into black-box attacks (MM-SafetyBench [27], SI-Attack [50], and FigStep [14]) and white-box attacks (HADES [23], VA-Jailbreak [31]).
@@
Figure 3: ASR under different prompts against MLLMs.
@@
In this section, we explore the performance of $\mathsf{FC\text{-}Attack}$ and conduct ablation study and defense research. We conduct jailbreak experiments on MLLMs for $\mathsf{FC\text{-}Attack}$. As shown in Figure A1 , it is a successful jailbreak case on Gemini-1.5.
Table 2: ASR comparison across models and attack shapes/sizes.
@@
Jailbreaking via Images. In Table 1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models. We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both models compared to all baselines. For example, the ASRs are $94\%$, $92\%$, $90\%$, and $90\%$ on Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, respectively. However, the ASR on some production models, such as Claude-3.5, GPT-4o, and GPT-4o mini, is relatively low, at $4\%$, $30\%$, and $10\%$, respectively. This might be because these production models have more advanced and updated visual safety alignment strategies.
@@
Jailbreaking via Videos. To conduct attacks from the video modality, we transform each jailbreak image into a 3-second video by setting all frames into the same image. Note that we also consider the Procedure Flowcharts, where each part (1 question and 5 steps) has been sequentially filled into a 0.5s video frame, resulting in a 3s video. We then evaluate the effectiveness of video jailbreak on three models: Qwen-vl-max, Qwen2.5-Omni and LLaVA-Video. The performance is summarized in Table 4. Our $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves a stable 88% ASR, whereas HADES peaks at 46% on Qwen-vl-max (dropping to 28% on LLaVA-Video) and Figstep fluctuates between 78% on Qwen-vl-max and 2% on Qwen2.5-Omni, highlighting our method’s consistent performance across models. As shown in Figure A2, jailbreaks using harmful text have an extremely low ASR. When the same harmful queries and steps are delivered via the video modality, the MLLMs become highly vulnerable, with ASR up to 88%.
Table 4: Comparison of ASR for different methods and models.
Method ASR (%)
Qwen-vl-max Qwen2.5-Omni LLaVA-Video
HADES 18 40 28
Figstep 78 2 10
Ours (Vertical) 72 58 76
Ours (Ensemble) 88 86 88
Ours (Procedure) 72 28 82